# CODIS: Benchmarking Context-Dependent Visual Comprehension for Multimodal Large Language Models

### v2

@@ Abstract
[…]
@@ 2.2 Visual Ambiguities
Visual ambiguities arises from absence of visual information or other noises (Denison et al., 2018). It is an important phenomenon in human visual cognition. Previous works have investigated visual ambiguities caused by optical illusions (Guan et al., 2023; Cui et al., 2023; Fu et al., 2023b). Images which may cause illusions are leveraged to evaluate visual hallucinations in MLLMs. Other works explore ambiguities in visual event classification (Sanders et al., 2022; Schmarje et al., 2022; Rajeswar et al., 2022), focusing on model robustness against ambiguous images. Different from these works, we assess capability of MLLMs to disambiguate images instead of just recognizing ambiguities.
@@ 3.1 Taxonomy of Context
Attributes. To comprehensively understand objects in a scene, it is not enough to solely rely on the visual appearances we see. Non-visible attributes of objects are also important. We consider two types of attributes, including physical attributes, such as structure, size, temperature, smell and touch, as Figure 2(g) shows, and abstract attributes, such as function and mechanism, as Figure 2(h) shows.
@@
mPLUG-Owl-2-7B (Ye et al., 2023) ~7B CLIPViT-L LLaMA-2-7B Attention
@@ 4.1 Models
As shown in Table 2, we perform evaluation on 14 popular MLLMs, which are divided into three groups: (1) API-based models: GPT-4V (OpenAI, 2023) and Gemini (Gemini Team et al., 2023); (2) Open-source ~7B models: mPLUG-Owl-2-7B (Ye et al., 2023), MiniGPT-4-7B (Zhu et al., 2023), LLaVA-1.5-7B (Liu et al., 2023a), InstructBLIP-7B (Dai et al., 2023), Otter-7B (Li et al., 2023a), LLaVA-7B (Liu et al., 2023b), Qwen-VL (Bai et al., 2023), OpenFlamingo (Awadalla et al., 2023) and BLIP-2-6.7B (Li et al., 2023d); (3) Open-source ~13B models: LLaVA-1.5-13B, BLIP-2-11B and InstructBLIP-13B.
@@ 4.2 Results
Overall results. From an overall perspective, human is far ahead of all the MLLMs in all the categories, indicating that MLLMs have tremendous potential of improvement on context-dependent visual comprehension. API-based models significantly outperform open-source models, and GPT-4V gets the highest score among all the MLLMs. In terms of open-source models, mPLUG-Owl-2-7B has the best performance on $\mathrm{Acc}_{p}$ and LLaVA-1.5-13B has the best performance on $\mathrm{Acc}_{q}$.
@@ 5.1 Visual Information Extraction
GPT-4V 14:12 07:12 16:9
Gemini 18:11 05:21 22:2
mPLUG-Owl-2-7B 23:6 09:15 23:1
LLaVA-1.5-13B 21:8 19:7 22:4
@@ Appendix E Detailed Results of GPT-4 Evaluation
I’ll give you an image and some additional context, which provides information closely related to the scene of the picture. Please answer my question based on the image and the context. Be sure to refer to the context and extract necessary information from it to help you answer the question because it contains helpful information that is not included in the image.
Your answer should contain two parts. Two parts should be separated by a newline. In the first part, please think of the question step by step based on the image and context and output your reasoning process. In the second part, please summarize your reasoning process and directly answer the question in a single word or phrase.
Context: [CONTEXT HERE]
Question: [QUESTION HERE]
@@
Evaluation
Please evaluate the output of models based on the given question and groundtruth and tell me whether the output is right.
Please pay attention to the following rules:
1. The output contains rationale of the reasoning process and answer which is summarized from the reasoning process. Please extract the answer from the output and make your judgement only based on answer, NOT rationale.
2. The answer is right if it follows the question in meaning and is consistent with the groundtruth.
3. Do not be too strict about the answer. Format different from the groundtruth and minor grammar issues are allowed.
If you think the answer is correct according to the groundtruth, please output "right", otherwise output "wrong". You can only print "right" or "wrong" and nothing else.
Here is the question: [QUESTION HERE]
Here is the groundtruth: [GROUNDTRUTH HERE]
Here is the output: [OUTPUT HERE]
@@
Appendix F Model Inference and Evaluation Prompts
We report prompt for model inference in Table 11 and prompt for GPT-4 evaluation in Table 12. The same prompt is employed for inference across all models.
Appendix G More Cases of CODIS
We present additional cases of CODIS, showcasing outputs of MLLMs on these cases in Figure 6, Figure 7, Figure 8, Figure 9 and Figure 10. We select two cases from each of the five categories. To aid readers in better comprehension, we provide explanations regarding the ambiguities in each image. It is important to note that these explanations are not visible to MLLMs or human volunteers who answered the questions in CODIS.

### v3

@@ Abstract
Fuwen Luo1∗, Chi Chen1∗, Zihao Wan1, Zhaolu Kang6, Qidong Yan5, Yingjie Li5, Xiaolong Wang1, Siyu Wang2, Ziyue Wang1, Xiaoyue Mi7, Peng Li2,3†, Ning Ma5, Maosong Sun1†, Yang Liu1,2,3,4 1Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2Institute for AI Industry Research (AIR), Tsinghua University, Beijing, China 3 Shanghai Artificial Intelligence Laboratory, Shanghai, China, 4 Jiangsu Collaborative Innovation Center for Language Competence, Jiangsu, China 5Key Laboratory of Linguistic and Cultural Computing Ministry of Education, Northwest Minzu University, China 6College of Software, Jilin University, China 7Institute of Computing Technology, Chinese Academy of Sciences {lfw23,chenchi19}@mails.tsinghua.edu.cn lipeng@air.tsinghua.edu.cn, sms@tsinghua.edu.cn
@@ 2.2 Visual Ambiguities
Visual ambiguities arises from absence of visual information or other noises (Denison et al., 2018). It is an important phenomenon in human visual cognition. Previous works investigate visual ambiguities caused by optical illusions (Guan et al., 2023; Cui et al., 2023; Fu et al., 2023b). Images which may cause illusions are leveraged to evaluate visual hallucinations in MLLMs. Other works explore ambiguities in visual event classification (Sanders et al., 2022; Schmarje et al., 2022; Rajeswar et al., 2022), focusing on model robustness against ambiguous images. Different from these works, we assess capability of MLLMs to disambiguate images instead of just recognizing ambiguities.
@@ 3.1 Taxonomy of Context
Attributes. To comprehensively understand objects in a scene, it is not enough to solely rely on visual appearances we see. Non-visible attributes are also important. We consider two types of attributes, including physical attributes, such as structure, size, temperature, smell and touch, as Figure 2(g) shows, and abstract attributes, such as function and mechanism, as Figure 2(h) shows.
@@
mPLUG-Owl-2 (Ye et al., 2023) ~7B CLIPViT-L LLaMA-2-7B Attention
@@ 4.1 Models
As shown in Table 2, we perform evaluation on 14 popular MLLMs, which are divided into three groups: (1) API-based models: GPT-4V (OpenAI, 2023) and Gemini (Gemini Team et al., 2023); (2) Open-source ~7B models: mPLUG-Owl-2 (Ye et al., 2023), MiniGPT-4-7B (Zhu et al., 2023), LLaVA-1.5-7B (Liu et al., 2023a), InstructBLIP-7B (Dai et al., 2023), Otter-7B (Li et al., 2023a), LLaVA-7B (Liu et al., 2023b), Qwen-VL (Bai et al., 2023), OpenFlamingo (Awadalla et al., 2023) and BLIP-2-6.7B (Li et al., 2023d); (3) Open-source ~13B models: LLaVA-1.5-13B, BLIP-2-11B and InstructBLIP-13B.
@@ 4.2 Results
Overall results. From an overall perspective, human is far ahead of all the MLLMs in all the categories, indicating that MLLMs have tremendous potential of improvement on context-dependent visual comprehension. API-based models significantly outperform open-source models, and GPT-4V gets the highest score among all the MLLMs. In terms of open-source models, mPLUG-Owl-2 has the best performance on $\mathrm{Acc}_{p}$ and LLaVA-1.5-13B has the best performance on $\mathrm{Acc}_{q}$.
@@ 5.1 Visual Information Extraction
GPT-4V 14:12 07:12 16:90
Gemini 18:11 05:21 22:20
mPLUG-Owl-2 23:60 09:15 23:10
LLaVA-1.5-13B 21:80 19:70 22:40
@@ 6 Conclusion
Acknowledgements
This work is supported by the National Key R&D Program of China (2022ZD0160502) and the National Natural Science Foundation of China (No. 61925601, 62276152). We thank Sijie Cheng, Zhicheng Guo for their advice in paper writing. We thank Zonghan Yang, Yile Wang and Yuanchi Zhang for their valuable discussions.
@@ Appendix E Detailed Results of GPT-4 Evaluation
w/ DI w/ CoT I’ll give you an image and some additional context, which provides information closely related to the scene of the picture. Please answer my question based on the image and the context. Be sure to refer to the context and extract necessary information from it to help you answer the question because it contains helpful information that is not included in the image. Your answer should contain two parts. Two parts should be separated by a newline. In the first part, please think of the question step by step based on the image and context and output your reasoning process. In the second part, please summarize your reasoning process and directly answer the question in a single word or phrase. Context: [CONTEXT HERE] Question: [QUESTION HERE]
w/ DI w/o CoT I’ll give you an image and some additional context, which provides information closely related to the scene of the picture. Please answer my question based on the image and the context. Be sure to refer to the context and extract necessary information from it to help you answer the question because it contains helpful information that is not included in the image. Please answer in a single word or phrase. Context: [CONTEXT HERE] Question: [QUESTION HERE]
w/o DI w/ CoT I’ll give you an image and some additional context, which provides information closely related to the scene of the picture. Please answer my question based on the image and the context. Your answer should contain two parts. Two parts should be seperated by a newline. In the first part, please think of the question step by step based on the image and context and output your reasoning process. In the second part, please summarize your reasoning process and directly answer the question in a single word or phrase. Context: [CONTEXT HERE] Question: [QUESTION HERE]
w/o DI w/o CoT I’ll give you an image and some additional context, which provides information closely related to the scene of the picture. Please answer my question based on the image and the context. Please answer in a single word or phrase. Context: [CONTEXT HERE] Question: [QUESTION HERE]
@@
Please evaluate the output of models based on the given question and groundtruth and tell me whether the output is right. Please pay attention to the following rules: 1. The output contains rationale of the reasoning process and answer which is summarized from the reasoning process. Please extract the answer from the output and make your judgement only based on answer, NOT rationale. 2. The answer is right if it follows the question in meaning and is consistent with the groundtruth. 3. Do not be too strict about the answer. Format different from the groundtruth and minor grammar issues are allowed. If you think the answer is correct according to the groundtruth, please output "right", otherwise output "wrong". You can only print "right" or "wrong" and nothing else. Here is the question: [QUESTION HERE] Here is the groundtruth: [GROUNDTRUTH HERE] Here is the output: [OUTPUT HERE]
@@
Model w/ DI w/ CoT w/ DI w/o CoT w/o DI w/ CoT w/o DI w/o CoT
GPT-4V 40.0 22.0 34.0 25.0
Gemini 32.0 16.0 27.0 11.0
mPLUG-Owl-2 13.0 14.0 11.0 13.0
LLaVA-1.5-13B 15.0 04.0 11.0 06.0
Table 13: Model performances with different prompts. We report $\mathrm{Acc}_{p}$.
Appendix F Prompt for Model Inference
As shown in Table 11, we design four prompts for model inference, and investigate impacts of detailed instructions and chain of thought prompting on model performance.
  • Detailed Instructions (DI): We stated that contexts contain helpful information that is not included in the image and encourage models to refer to contexts.
  • Chain of Thought Prompting (CoT): We asked models to think of the questions step by step, and then summarize the reasoning processes and answer the question in single words or phrases.
We randomly select 200 pairs, 400 queries from our benchmark and report $\mathrm{Acc}_{p}$ scores of GPT-4V, Gemini, mPLUG-Owl-2 and LLaVA-1.5-13B with the prompts above. The results are shown in Table 13.
The results indicated that DI and CoT help improve model performances. There were no significant differences in performance observed for LLaVA-1.5-13B when using either of the prompts, while the other models exhibit their best performances with DI and CoT. Therefore, based on the results, we decided to use the prompt with both DI and CoT for all the models in our evaluation process.
Additionally, it is worth noting that there are subtle variations in input formats and settings of different models, including placement of images, use of special tokens, etc. Throughout our evaluation, we adhere to the default configurations of each model. Details regarding input formats and settings of our evaluated MLLMs can be found in their documentations or repositories.
Appendix G Prompt for GPT-4 Evaluation
We report prompt for GPT-4 evaluation in Table 12. The results of GPT-4 evaluation in Table 7 and results in Table 10 are both based on the prompt.
Appendix H More Cases of CODIS
We present additional cases of CODIS, showcasing outputs of MLLMs on these cases in Figure 6, Figure 7, Figure 8, Figure 9 and Figure 10. We select two cases from each of the five categories. The incorrect model responses are highlighted in red. To aid readers in better comprehension, we provide explanations regarding the ambiguities in each image. It is important to note that these explanations are not visible to MLLMs or human volunteers who answered the questions in CODIS.